OpenStudy (anonymous):

will give medal and fan 2x+y=-10 explain (-2,-6) (4,-18) and (-1,8)

OpenStudy (anonymous):

What are you asked to do?

OpenStudy (anonymous):

solve them to see which has the right solution

OpenStudy (anonymous):

Plug each point in 2x+y=-10 and check if it's a true statement. I'll do the first one as an example: (-2,-6) 2(-2)+(-6)=-10 -4-6=-10 -10=-10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

Alright, but open a new question so more people can see it and help .